# Online Resource Allocation via Static Bundle Pricing

### v1

@@ Appendix F Tree Setting
For every edge $e$ and buyer $b$ it holds that:
$\mathbb{E}_{M,I}\left[\ell_{\textnormal{{UAlg}}_{M,I}\setminus\{b\}}(e)\mid b\in\textnormal{{UAlg}}_{M,I}\right]\leq\mathbb{E}_{M,I}\left[\ell_{\textnormal{{UAlg}}_{M,I}}(e)\right]$
@@
Fix an edge $e$, a buyer $b$ and the set $\mathcal{A}$ of buyers that desire the same set of edges, $S_{b}$. Define also $\mathcal{A}_{M,I}\coloneqq\textnormal{{UAlg}}_{M,I}\cap\mathcal{A}$. It suffices to prove that:
$\mathbb{E}_{M,I}\left[\ell_{\mathcal{A}_{M,I}\setminus\{b\}}(e)\mid b\in\textnormal{{UAlg}}_{M,I}\right]\leq\mathbb{E}_{M,I}\left[\ell_{\mathcal{A}_{M,I}}(e)\right]$
since whether $b\in\textnormal{{UAlg}}_{M,I}$ does not affect buyers that desire sets different from $S_{b}$ (as we will formalize when we establish subadditivity of the menu).
First, if we offer $|\mathcal{A}|$ copies of bundle $S_{b}$ then for every buyer $b^{\prime}\in\mathcal{A}$ the event $b^{\prime}\in\mathcal{A}_{M,I}$ depends only on whether buyer $b^{\prime}$ is realized thus the lemma trivially holds. Otherwise, one copy of bundle $S_{b}$ is offered with some probability. Since $b\in\mathcal{A}_{M,I}$, buyer $b$ got the single copy of bundle $S_{b}$, thus $\ell_{\mathcal{A}_{M,I}\setminus\{b\}}(e)=0$, proving the lemma. ∎
We now prove the following two lemmas.
Lemma F.3.
$\mathbb{E}_{M,I}\left[\textsc{NonPeak}_{M,I}\right]\leq 0.1\,\mathbb{E}_{M,I}\left[\textnormal{{UAlg}}_{M,I}\right]$
@@
It holds that:
[…]
$\leq\sum_{b\in\textnormal{{UAlg}}_{M,I}}\ell_{\textnormal{{UAlg}}_{M,I}\setminus\{b\}}(f_{b})$
$\leq\sum_{b\in\mathcal{N}}\mathbbm{1}{\left\{b\in\textnormal{{UAlg}}_{M,I}\right\}}\ell_{\textnormal{{UAlg}}_{M,I}\setminus\{b\}}(f_{b})$
Taking expectations in both parts we get:
$\mathbb{E}_{M,I}\left[\textsc{NonPeak}_{M,I}\right]$ $\leq\sum_{b\in\mathcal{N}}\mathbb{P}\!\left[b\in\textnormal{{UAlg}}_{M,I}\right]\mathbb{E}_{M,I}\left[\ell_{\textnormal{{UAlg}}_{M,I}\setminus\{b\}}(f_{b})\mid b\in\textnormal{{UAlg}}_{M,I}\right]$
$\leq\sum_{b\in\mathcal{N}}\mathbb{P}\!\left[b\in\textnormal{{UAlg}}_{M,I}\right]\mathbb{E}_{M,I}\left[\ell_{\textnormal{{UAlg}}_{M,I}}(f_{b})\right]$
$=0.1\cdot\mathbb{E}_{M,I}\left[\textnormal{{UAlg}}_{M,I}\right],$
where in the second inequality we used Lemma˜F.2. ∎
Lemma F.4.
$\mathbb{E}_{M,I}\left[\textsc{Peak}_{M,I}\right]\leq 0.2\;\mathbb{E}_{M,I}\left[\textnormal{{UAlg}}_{M,I}\right]$
Proof.
It holds that:
[…]
$\leq\sum_{b\in\textnormal{{UAlg}}_{M,I}}\mathbbm{1}{\left\{\ell_{\textnormal{{Alg}}_{M,I}}(e_{b}^{1})>0\right\}}+\mathbbm{1}{\left\{\ell_{\textnormal{{Alg}}_{M,I}}(e_{b}^{2})>0\right\}}$
$\leq\sum_{b\in\mathcal{N}}\mathbbm{1}{\left\{b\in\textnormal{{UAlg}}_{M,I}\right\}}\left(\ell_{\textnormal{{UAlg}}_{M,I}\setminus\{b\}}(e_{b}^{1})+\ell_{\textnormal{{UAlg}}_{M,I}\setminus\{b\}}(e_{b}^{2})\right)$
Taking expectations in both parts we get similarly to the non-peak case:
$\mathbb{E}_{M,I}\left[\textsc{Peak}_{M,I}\right]$ $\leq\sum_{b\in\mathcal{N}}\mathbb{P}_{M,I}\!\left[b\in\textnormal{{UAlg}}_{M,I}\right]\left(\mathbb{E}_{M,I}\left[\ell_{\textnormal{{UAlg}}_{M,I}}(e_{b}^{1})\right]+\mathbb{E}_{M,I}\left[\ell_{\textnormal{{UAlg}}_{M,I}}(e_{b}^{2})\right]\right)$
$\leq\sum_{b\in\mathcal{N}}\mathbb{P}_{M,I}\!\left[b\in\textnormal{{UAlg}}_{M,I}\right]\left(0.1+0.1\right)$
$=0.2\;\mathbb{E}_{M,I}\left[\textnormal{{UAlg}}_{M,I}\right]$
From Corollary˜F.1, Lemma˜F.4, Lemma˜F.3 and taking expectations in Equation˜2 we get:
$\mathbb{E}_{M,I}\left[\textnormal{{Alg}}_{M,I}\right]\geq 0.7\,\mathbb{E}_{M,I}\left[\textnormal{{UAlg}}_{M,I}\right]\geq\frac{7}{800}\textnormal{{FracOpt}}$
Appendix G Deferred Proofs
G.1 Proof of Lemma 2.1
Proof.
Consider the optimal allocation for a fixed realization of buyers’ values. For each buyer $b$ and value $v\in\mathcal{S}$, let $y_{b,v}\in\{0,1\}$ indicate whether in this realization buyer $b$ is allocated bundle $S_{b}$ when their value is $v$. By feasibility of the offline solution, for every $e\in\mathcal{M}$ we have
$\sum_{b:e\in S_{b}}\sum_{v\in\mathcal{S}}y_{b,v}\leq c(e),$
$\sum_{b:e\in S_{b}}\sum_{v\in\mathcal{S}}x_{b,v}=\mathbb{E}_{I}\left[\sum_{b:e\in S_{b}}\sum_{v\in\mathcal{S}}y_{b,v}\right]\leq c(e),$
$\sum_{b\in[n]}\sum_{v\in\mathcal{S}}x_{b,v}\cdot v=\mathbb{E}_{I}\left[\sum_{b\in[n]}\sum_{v\in\mathcal{S}}y_{b,v}\cdot v\right]=\textnormal{{Opt}},$
G.2 Proof of Lemma 2.3
Proof.
Fix a buyer $b$ and consider the vector $X^{(b)}=(x_{b,v})_{v\in\mathcal{S}}$, where values in $\mathcal{S}$ are ordered increasingly. Recall the value‐exchange argument from the main text: whenever $v^{\prime}>v$, we may shift an infinitesimal amount of mass from $x_{b,v}$ to $x_{b,v^{\prime}}$ without violating any LP constraints, unless one of the box constraints $0\leq x_{b,v}$ or $x_{b,v^{\prime}}\leq q_{b,v^{\prime}}$ becomes tight. Such a shift strictly increases the objective because $v^{\prime}>v$.
@@ Appendix G Deferred Proofs
Fix a bundle $S$ and let $\mathcal{A}=\mathcal{A}_{S}$ and $X=X^{(S)}$ for simplicity. Recall that
$x_{S,v}\;=\;\sum_{b\in\mathcal{A}}x_{b,v}\qquad\text{and}\qquad q_{S,v}\;=\;\sum_{b\in\mathcal{A}}q_{b,v},$
@@
Lemma G.1.
If any buyer $b\in\mathcal{A}$ has a crucial value, then all crucial values across buyers in $\mathcal{A}$ occur at the same value. We call this value the crucial value of $X$.
@@
Suppose buyer $b$ has crucial value $w_{b}$ and buyer $b^{\prime}$ has crucial value $w_{b^{\prime}}$ with $w_{b}<w_{b^{\prime}}$. Because both buyers demand the same bundle $S$, shifting a small amount of mass from $x_{b,w_{b}}$ to $x_{b^{\prime},w_{b^{\prime}}}$ preserves all LP capacity constraints: both variables consume exactly the same set of items. This shift increases the objective because $w_{b^{\prime}}>w_{b}$, contradicting optimality of the LP solution. Hence no two buyers in $\mathcal{A}$ can have crucial values at different levels, proving the claim. ∎
Lemma G.2.
@@
Since $w$ is crucial for $X$, we have $0<x_{S,w}<q_{S,w}$. First suppose there exists $v<w$ with $x_{S,v}>0$. Then there is some buyer $b\in\mathcal{A}$ with $x_{b,v}>0$. Because $x_{S,w}<q_{S,w}$, there exists some buyer $b^{\prime}\in\mathcal{A}$ with $x_{b^{\prime},w}<q_{b^{\prime},w}$.
@@
Now suppose there exists $u>w$ with $x_{S,u}<q_{S,u}$. Then there is a buyer $b^{\prime}\in\mathcal{A}$ with $x_{b^{\prime},u}<q_{b^{\prime},u}$. Since $x_{S,w}>0$, there exists a buyer $b\in\mathcal{A}$ with $x_{b,w}>0$. Again, applying the value–exchange argument, we can shift an infinitesimal amount of mass from $x_{b,w}$ to $x_{b^{\prime},u}$. Feasibility is preserved (both entries use the same bundle $S$), and the objective strictly increases because $u>w$, contradicting optimality. Thus no such $u$ can exist, and we must have $x_{S,u}=q_{S,u}$ for all $u>w$. This proves the claim. ∎
We now conclude the proof of Lemma˜2.4. If no crucial value exists in $X$, then for every $v$ we have either $x_{S,v}=0$ or $x_{S,v}=q_{S,v}$, and by the same exchange argument as above, zeros must occur only at lower values and tight values only at higher values. If a crucial value exists, then by Lemma˜G.1 it is unique, and by Lemma˜G.2 $X$ consists of zeros below it, the crucial value, and tight entries above it. Thus in all cases the vector $X$ has exactly the form stated in the lemma. ∎
G.4 Proof of Lemma 7.3
@@
$g(m,t,B+1)\;\geq\;\frac{e(B+1)}{t}\exp\!\left(\frac{m}{(B+1)t^{B+1}}\right).$
@@
therefore
@@
$g(m,t,B+1)\;\geq\;\frac{e(B+1)}{t}\exp\!\left(\frac{2Bt\ln m}{B+1}\right)\;=\;\frac{e(B+1)}{t}\,m^{2Bt/(B+1)}.$
@@
$\frac{e(B+1)}{t}\,m^{2Bt/(B+1)}\;\geq\;Bt^{t}.$
Taking logarithms and using $t\leq m^{1/(B+2)}$ (and $m$ sufficiently large) shows that the left-hand side grows much faster than the right-hand side as a function of $m$. Hence the inequality holds, and
$g(m,t,B+1)\;\geq\;Bt^{t},$
as claimed. ∎
G.5 Proof of Lemma 7.7
Proof.
Apply Lemma˜7.6 with $n=m=dt$, $r=B+1$, and $d=n/t$:
$g_{\mathrm{bal}}(m,t,B+1,d)\;\geq\;\frac{e(B+1)}{t}\cdot\exp\!\left(\frac{m}{(B+1)t^{B+1}}\right)=\frac{e(B+1)}{t}\cdot\exp\!\left(\frac{d}{(B+1)t^{B}}\right).$
By the definition of $t$,
Hence
$g_{\mathrm{bal}}(m,t,B+1,d)\;\geq\;d^{\Omega(t)}.$
G.6 Auxiliary Lemma
Lemma G.3.
@@ G.6 Auxiliary Lemma
It is a standard fact for Poisson binomial random variables (i.e., sums of independent Bernoulli variables) that any median $m$ of $X$ satisfies $\lfloor\mu\rfloor\leq m\leq\lceil\mu\rceil$ (see for example Theorem 3.2 and Corollary 3.1 in [jogdeo]). In particular, $m\geq\lfloor\mu\rfloor$. Since $\mu\geq 1$ implies $\lfloor\mu\rfloor\geq\mu/2$, we obtain

### v2

@@ F.1 Fractional Relaxation and Offline Optimum
$\textnormal{{FracOpt}}\geq\textnormal{{Opt}}$.
@@
Consider the optimal allocation for a fixed realization of buyers’ values. For each buyer $b$ and value $v\in\mathcal{U}$, let $y_{b,v}\in\{0,1\}$ indicate whether in this realization buyer $b$ is allocated bundle $S_{b}$ when their value is $v$. By feasibility of the offline solution, for every $e\in\mathcal{M}$ we have
$\sum_{b:e\in S_{b}}\sum_{v\in\mathcal{U}}y_{b,v}\leq c(e),$
$\sum_{b:e\in S_{b}}\sum_{v\in\mathcal{U}}x_{b,v}=\mathbb{E}_{I}\left[\sum_{b:e\in S_{b}}\sum_{v\in\mathcal{U}}y_{b,v}\right]\leq c(e),$
$\sum_{b\in[n]}\sum_{v\in\mathcal{U}}x_{b,v}\cdot v=\mathbb{E}_{I}\left[\sum_{b\in[n]}\sum_{v\in\mathcal{U}}y_{b,v}\cdot v\right]=\textnormal{{Opt}},$
F.2 Proof of Lemma 5.1
@@
Fix a buyer $b$ and consider the vector $X^{(b)}=(x_{b,v})_{v\in\mathcal{U}}$, where values in $\mathcal{U}$ are ordered increasingly. Recall the value‐exchange argument from the main text: whenever $v^{\prime}>v$, we may shift an infinitesimal amount of mass from $x_{b,v}$ to $x_{b,v^{\prime}}$ without violating any LP constraints, unless one of the box constraints $0\leq x_{b,v}$ or $x_{b,v^{\prime}}\leq q_{b,v^{\prime}}$ becomes tight. Such a shift strictly increases the objective because $v^{\prime}>v$.
@@
Fix a bundle $S$ and let $X=X^{(S)}$ for simplicity. Recall that
$x_{S,v}\;=\;\sum_{b\in\mathcal{N}_{S}}x_{b,v}\qquad\text{and}\qquad q_{S,v}\;=\;\sum_{b\in\mathcal{N}_{S}}q_{b,v},$
@@
Lemma F.2.
If any buyer $b\in\mathcal{N}_{S}$ has a crucial value, then all crucial values across buyers in $\mathcal{N}_{S}$ occur at the same value. We call this value the crucial value of $X$.
@@
Suppose buyer $b$ has crucial value $w_{b}$ and buyer $b^{\prime}$ has crucial value $w_{b^{\prime}}$ with $w_{b}<w_{b^{\prime}}$. Because both buyers demand the same bundle $S$, shifting a small amount of mass from $x_{b,w_{b}}$ to $x_{b^{\prime},w_{b^{\prime}}}$ preserves all LP capacity constraints: both variables consume exactly the same set of items. This shift increases the objective because $w_{b^{\prime}}>w_{b}$, contradicting optimality of the LP solution. Hence no two buyers in $\mathcal{N}_{S}$ can have crucial values at different levels, proving the claim. ∎
Lemma F.3.
@@
Since $w$ is crucial for $X$, we have $0<x_{S,w}<q_{S,w}$. First suppose there exists $v<w$ with $x_{S,v}>0$. Then there is some buyer $b\in\mathcal{N}_{S}$ with $x_{b,v}>0$. Because $x_{S,w}<q_{S,w}$, there exists some buyer $b^{\prime}\in\mathcal{N}_{S}$ with $x_{b^{\prime},w}<q_{b^{\prime},w}$.
@@
Now suppose there exists $u>w$ with $x_{S,u}<q_{S,u}$. Then there is a buyer $b^{\prime}\in\mathcal{N}_{S}$ with $x_{b^{\prime},u}<q_{b^{\prime},u}$. Since $x_{S,w}>0$, there exists a buyer $b\in\mathcal{N}_{S}$ with $x_{b,w}>0$. Again, applying the value–exchange argument, we can shift an infinitesimal amount of mass from $x_{b,w}$ to $x_{b^{\prime},u}$. Feasibility is preserved (both entries use the same bundle $S$), and the objective strictly increases because $u>w$, contradicting optimality. Thus no such $u$ can exist, and we must have $x_{S,u}=q_{S,u}$ for all $u>w$. This proves the claim. ∎
We now conclude the proof of Lemma 5.2. If no crucial value exists in $X$, then for every $v$ we have either $x_{S,v}=0$ or $x_{S,v}=q_{S,v}$, and by the same exchange argument as above, zeros must occur only at lower values and tight values only at higher values. If a crucial value exists, then by Lemma F.2 it is unique, and by Lemma F.3 $X$ consists of zeros below it, the crucial value, and tight entries above it. Thus in all cases the vector $X$ has exactly the form stated in the lemma. ∎
F.4 Proof of Lemma 8.1
Proof.
A buyer is rejected only if some item in her bundle is oversold. Thus
$|\textsc{Blocked}(M,I)|\;\leq\;\sum_{e\in\mathcal{M}}\sum_{k=c(e)+1}^{\infty}\mathbbm{1}{\left\{L_{M,I}(e)=k\right\}}\,(k-c(e)).$
Taking expectations and using linearity,
$\mathbb{E}_{M,I}\left[|\textsc{Blocked}(M,I)|\right]\;\leq\;\sum_{e\in\mathcal{M}}\sum_{k=c(e)+1}^{\infty}(k-c(e))\,\mathbb{P}\!\left[L_{M,I}(e)=k\right].$
By Lemma 6.3, the random load $Z_{M,I}(e)$ is the sum of negatively associated Bernoulli variables with total mean $\mu_{e}\leq c(e)/\gamma$. A standard Chernoff bound for negatively associated variables gives
Summing over $j$ gives
$\sum_{k=c(e)+1}^{\infty}(k-c(e))\,\mathbb{P}\!\left[L_{M,I}(e)=k\right]\;\leq\;\frac{1}{20m}.$
Summing over all $m$ items yields the bound. ∎
F.5 Proof of Lemma 10.3
@@
$g(m,t,B+1)\;\geq\;\frac{(B+1)}{et}\exp\!\left(\frac{m}{(B+1)t^{B+1}}\right).$
@@
Hence
$t^{B+1}\;\leq\;\frac{m}{2B\ln m}\cdot\frac{1}{t},$
and therefore
@@
$g(m,t,B+1)\;\geq\;\frac{(B+1)}{et}\exp\!\left(\frac{2Bt\ln m}{B+1}\right)\;=\;\frac{(B+1)}{et}\,m^{2Bt/(B+1)}.$
@@
$\frac{(B+1)}{et}\,m^{\frac{2Bt}{B+1}}\;\geq\;Bt^{t}.$
Since we already established that $t\leq m^{1/(B+2)}$, we have $m\geq t^{B+2}$ and therefore
$\frac{(B+1)}{et}\,m^{\frac{2Bt}{B+1}}$ $\geq\frac{(B+1)}{et}\,\bigl(t^{B+2}\bigr)^{\frac{2Bt}{B+1}}=\frac{(B+1)}{et}\,t^{\frac{2B(B+2)}{B+1}\,t}$
$=Bt^{t}\cdot\frac{(B+1)}{eB}\,t^{\left(\frac{2B(B+2)}{B+1}-1\right)t-1}.$
Using $\left(\frac{2B(B+2)}{B+1}-1\right)t-1\geq 1$ (as $t\geq 1$), we get $g(m,t,B+1)\geq Bt^{t},$ as claimed. ∎
F.6 Auxiliary Lemma
Lemma F.4.
@@ F.6 Auxiliary Lemma
It is a standard fact for Poisson binomial random variables (i.e., sums of independent Bernoulli variables) that any median $m$ of $X$ satisfies $\lfloor\mu\rfloor\leq m\leq\lceil\mu\rceil$. In particular, $m\geq\lfloor\mu\rfloor$. Since $\mu\geq 1$ implies $\lfloor\mu\rfloor\geq\mu/2$, we obtain